OpenStudy (anonymous):

find the centre and radius of the circle given by the equation x^(2)+y^(2)-4x+6y-12=0.

OpenStudy (anonymous):

centre => (4/2,-6/2)=(2,-3) radius= sqrt(2^2+(-3)^2-(-12))=sqrt(4+9+12)=sqrt(25)=5

OpenStudy (anonymous):